# DFAMS: Dynamic-flow guided Federated Alignment based Multi-prototype Search

### v1

@@ 3.3 Multi-Prototype Knowledge Alignment & Routing
To bridge internal DIF representations and external distributed knowledge structure (C2), we map $\mathbf{z}$ to a semantic space using a projection $g_{\text{align}}$, producing $\mathbf{r}=g_{\text{align}}(\mathbf{z})$. We train this space with two contrastive stages and use it for prototype-guided routing.
Inter-KB Alignment. We apply supervised contrastive learning to model the boundary between different knowledge bases:
@@
Adaptive triggering. If the maximum similarity falls below a threshold $\tau$, i.e., $\max_{i}s_{i}<\tau$, the system abstains from retrieval.
@@
Semantic Routing. Otherwise, it identifies the top-$N$ most similar prototypes $\mathcal{I}$ and allocates a total of $T$ retrieval slots across knowledge bases. The number of documents assigned to each knowledge base $k$ is computed as:
$w_{k}=\left\lfloor\frac{\sum_{i\in\mathcal{I},\,k_{i}=k}s_{i}}{\sum_{k^{\prime}}\sum_{i\in\mathcal{I},\,k_{i}=k^{\prime}}s_{i}}\cdot T\right\rfloor.$ (9)
This two-step mechanism ensures that retrieval is triggered only when necessary and distributes resources based on prototype-level semantic relevance.
@@
Table 1: Performance comparison (%) on Wiki, Med, PEP, MMLU, and MIRAGE, where bold indicates the best result, and symbol slash ”/” denotes inapplicable metrics due to lack of retrieval (NoRAG) or single-source setup (Merged RAG).
@@ 4 Experiments
We conduct extensive experiments across multiple datasets to evaluate the effectiveness of DFAMS in Federated Retrieval settings. Our experiment is organized to answer the following key research questions:
@@
  • RQ2 (Section 4.3): How do the proposed component contribute to the observed performance improvements?
@@ 4.1 Experimental Setup
We implement DFAMS on four open-weight, decoder-only LLMs with varying scales to evaluate scalability and generalization: Qwen2.5-0.5B, Qwen2.5-3B, Qwen2.5-7B (Yang et al. 2024), and Llama3.1-8B (Grattafiori et al. 2024).
@@
DFAMS utilizes FAISS (Douze et al. 2024) for dense vector retrieval across three federated retrieval (FR) scenarios, each comprising multiple knowledge bases: Wiki-KB (10 semantically clustered segments from Wikipedia (Wu et al. 2025)), Med-KB (4 sources: PubMed, StatPearls, textbooks, and medical Wikipedia (Zhao et al. 2025)), and PEP-KB (4 collections of internal enterprise policy documents). For each query, the top-10 chunks are retrieved collectively from the selected sources across all relevant knowledge bases within the same scenario. Further details on the retrieval pipeline and indexing strategies are provided in Appendix D.
@@
Based on the above knowledge base configurations, we construct three in-domain evaluation benchmarks—Wiki, Med, and PEP—each aligned with one of the FR scenarios and following the data setup described in Section 3.1. To further evaluate out-of-domain (OOD) generalization, we test models trained on Wiki and Med using a filtered subset of MMLU (Hendrycks et al. 2020) and a subset of MIRAGE (Xiong et al. 2024), respectively, with retrieval grounded in their corresponding knowledge bases. Further construction details are provided in Appendix E.
@@
We compare DFAMS with several representative retrieval and routing strategies: No-RAG, Merged-RAG(merges all corpora into one retrieval space), and two prompt-based methods—Prompt and CoTPrompt (Wei et al. 2023). We also include two recent multi-source retrieval approaches: RAGRoute (Guerraoui et al. 2025) that employs a binary classifier for each knowledge base to select the Top-$k$ sources, and RopMura (Wu et al. 2025), a prototype-based multi-agent routing method. Implementation details are provided in Appendix F.
@@
We report three complementary metrics: (1) Cls Acc, which measures whether the predicted knowledge base(s) match the ground truth, including correct Others predictions for no-retrieval cases; (2) Recall, computed over Top-10 retrieved documents for retrieval-triggering queries, measuring the proportion of gold documents successfully retrieved excluding no-retrieval cases; (3) QA, the end-to-end answer accuracy, using accuracy(QA ACC) for multiple-choice and LLM-based scoring (QA Score) (Zheng et al. 2023) for open-ended questions. More metric details are available in Appendix G.
@@ 4.2 Main Result Analysis (RQ1)
Table 1 summarizes the performance of DFAMS compared with a range of baseline methods across three key metrics: Cls Acc, Recall, and QA Acc/Score. DFAMS consistently outperforms all baselines on both in-domain and out-of-domain (OOD) datasets.
@@
Compared with advanced multi-source retrieval methods, DFAMS consistently achieves higher CLS ACC and Recall across all datasets and model backbones. For instance, on Wiki with Qwen2.5-7B, DFAMS outperforms RAGRoute by +8.96% in Cls Acc (85.03% vs. 76.07%) and +3.79% in Recall (53.83% vs. 50.04%). These improvements stem from DFAMS’s fine-grained modeling of FID, enabling more accurate identification of query intent and relevant knowledge sources. This leads to better routing and more focused retrieval with less cross-domain noise. As a result, DFAMS consistently achieves the highest QA accuracy across all datasets. For example, on the Qwen2.5-7B backbone, it reaches 82.82% QA accuracy on Med and 78.94% on Wiki, outperforming both RopMura and RAGRoute.
@@ 4.3 Ablation Study (RQ2)
Dynamic Information Flow Modeling. We conduct ablation studies to test our central hypothesis: that DIF signals not only exist but can be effectively detected and exploited to improve model performance. We design two experimental settings: (1) Frozen LLM w/o Align-MLP: We remove the trainable Align-MLP and directly utilize the extracted FID. This setting investigates whether native LLM activations inherently encode subdomain-aware signals—i.e., whether meaningful associations between user queries and knowledge subdomains can be inferred without explicit alignment. (2) Full DFAMS (w/ Align-MLP): We enable the trainable Align-MLP to utilize DIF to assesse whether modeling the knowledge base on FID leads to improved alignment and enhanced downstream performance.
@@
We conduct experiments on both the Wiki and Med datasets. Table 4 shows that in setting (1), DFAMS (2000-dim) outperforms both random (2000-dim) and full-layer (3584-dim) baselines, achieving +14.6% accuracy gain on Wiki and +11.78% on Med over the random baseline. These results validate our hypothesis: FID captures query-subdomain associations and can be directly leveraged, even without further training. In setting (2), with Align-MLP enabled, DFAMS achieves +3.54% and +1.66% higher accuracy than the random and full-layer baselines on Wiki, and +5.34% and +2.16% on Med, respectively. These results highlight the benefit of leveraging DIF for knowledge base modeling, resulting in better alignment and improved downstream performance. Additional results and analyses are provided in Appendix D.
@@ 4.4 Sensitivity Analysis (RQ3)
As shown in Figure 3 (right), when top-$T$ is 3, it achieves the best balance (84.2% accuracy, 53.8% recall). Lower $T$ values miss relevant sources, while higher $T$ dilutes document allocation across knowledge bases, reducing the chance of retrieving key information.
5 Conclusions and Future Works
We propose DFAMS, a novel FR framework that explicitly models the DIF within LLMs to enhance query understanding and cross-source knowledge alignment. By leveraging gradient-based neuron attribution and Shapley value estimation, DFAMS identifies latent neural activation paths that reflect user intent and subdomain relevance.The framework also introduces a Multi-Prototype Knowledge Alignment and Routing strategy, which enables fine-grained modeling of individual knowledge bases, effectively preserving inter-source semantic associations while supporting multi-knowledge-base routing. Our experiments show that DFAMS consistently outperforms existing FR baselines in knowledge base classification, retrieval quality, and end-to-end QA performance, confirming its effectiveness in resolving semantic ambiguities and improving cross-source routing in complex settings.
@@ 5 Conclusions and Future Works
Appendix
A. Notations Table
@@ II Number of isolated knowledge bases (KBs)
$\mathcal{D}={\mathcal{K}_{1},\dots,\mathcal{K}_{I}}$ KBs representing non-parameterized knowledge
$R$ Retrieved subset of KBs for answering the query
@@
$y$ Ground-truth answer (or response) to query $x$
@@
$g_{\text{align}}$ Projection function (Align MLP) mapping $\mathbf{z}$ to aligned query embedding $\mathbf{r}$ in the semantic space
@@ B. Algorithm
16:Select top-$T$ layers with highest total Shapley mass
17:Select top neuron groups of size $G$ per chosen layer.
18:return $\mathcal{N}$ using Equation 5
@@
3:Step 0: DIF Embedding Extraction
@@
C. Prompt

### v2

@@ 3.3 Multi-Prototype Knowledge Alignment & Routing
To bridge internal DIF representations and external distributed knowledge structure (C2), we map $\mathbf{z}$ to a semantic space using a projection $g_{\text{align}}$, producing $\mathbf{r}=g_{\text{align}}(\mathbf{z})$. We train it with two contrastive stages and use it for prototype-guided routing.
Inter-KB Alignment.
We apply supervised contrastive learning to model the boundary between different knowledge bases:
@@
Adaptive Triggering.
If the maximum similarity falls below a threshold $\tau$, i.e., $\max_{i}s_{i}<\tau$, the system abstains from retrieval.
@@
Semantic Routing.
Otherwise, it identifies the top-$N$ most similar prototypes $\mathcal{I}$ and allocates a total of $T$ retrieval slots across knowledge bases. The number of documents assigned to each knowledge base $k$ is computed as:
$w_{k}=\left\lfloor\frac{\sum_{i\in\mathcal{I},\,k_{i}=k}s_{i}}{\sum_{k^{\prime}}\sum_{i\in\mathcal{I},\,k_{i}=k^{\prime}}s_{i}}\cdot T\right\rfloor$ (5)
This two-step mechanism ensures retrieval occurs only when necessary, with resources allocated by prototype-level semantic relevance.
@@
Table 1: Performance comparison (%) on Wiki, Med, PEP, MMLU, and MIRAGE, where bold indicates the best result, and symbol slash "/" denotes inapplicable metrics .
@@ 4 Experiments
We conduct extensive experiments across multiple datasets to evaluate the effectiveness of DFAMS in Federated Retrieval settings and answer the following key research questions:
@@
  • RQ2 (Section 4.3): How do the proposed component contribute to performance improvements?
@@ 4.1 Experimental Setup
We implement DFAMS on four open-weight LLMs with varying scales to evaluate scalability and generalization: Qwen2.5-0.5B, Qwen2.5-3B, Qwen2.5-7B Yang et al. (2024), and Llama3.1-8B Grattafiori et al. (2024).
@@
Following Wu et al. (2025) and Guerraoui et al. (2025), DFAMS utilizes FAISS Douze et al. (2024) for dense vector retrieval across three FR scenarios, each comprising multiple knowledge bases. For each query, top-$10$ chunks are retrieved from selected sources across all relevant knowledge bases within the same scenario. Further details are provided in Appendix C
@@
We construct three in-domain evaluation benchmarks, Wiki, Med, and PEP. To further evaluate out-of-domain generalization, we test models trained on Wiki and Med using a subset of MMLU Hendrycks et al. (2020) and a subset of MIRAGE Xiong et al. (2024), following the setup of  Guerraoui et al. (2025) . Further construction details are provided in Appendix D.
@@
We compare DFAMS with several representative retrieval and routing strategies: No-RAG, Merged-RAG , and two prompt-based methods Prompt and CoT Prompt Wei et al. (2023). We further compare with a supervised fine-tuning baseline (SFT) Hu et al. (2021). In addition, we consider two recent multi-source retrieval approaches: RAGRoute Guerraoui et al. (2025), which employs a binary classifier for each knowledge base to select the Top-$k$ sources, and RopMura Wu et al. (2025), a prototype-based multi-agent routing method. Implementation details are provided in Appendix E.
@@
We report three complementary metrics: (1) Cls Acc, which measures whether the predicted knowledge base(s) match the ground truth, including correct Others predictions for no-retrieval cases; (2) Recall, computed over Top-10 retrieved documents for retrieval-triggering queries, measuring the proportion of gold documents successfully retrieved excluding no-retrieval cases; (3) QA, the end-to-end answer accuracy, using accuracy(QA ACC) for multiple-choice and LLM-based scoring (QA Score) Zheng et al. (2023) for open-ended questions. More metric details are available in Appendix F.
@@ 4.2 Main Result Analysis (RQ1)
Table 1 summarizes the performance of DFAMS compared with a range of baseline methods across three key metrics: Cls Acc, Recall, and QA Acc/Score. DFAMS consistently outperforms all baselines on both in-domain and out-of-domain datasets.
@@
Compared with advanced multi-source retrieval methods, DFAMS consistently achieves higher Cls Acc and Recall across all datasets and backbones. For instance, on Wiki with Qwen2.5-7B, DFAMS outperforms RAGRoute by +8.96% in Cls Acc (85.03% vs. 76.07%) and +3.79% in Recall (53.83% vs. 50.04%). These improvements stem from DFAMS’s fine-grained modeling of DIF, enabling more accurate identification of query intent and relevant knowledge sources. This leads to better routing and more focused retrieval with less cross-domain noise. As a result, DFAMS consistently achieves the highest QA accuracy across all datasets. For example, on the Qwen2.5-7B backbone, it reaches 82.82% QA accuracy on Med and 78.94% on Wiki, outperforming both RopMura and RAGRoute.
@@ 4.3 Ablation Study (RQ2)
Dynamic Information Flow Modeling. We conduct ablation studies to test our central hypothesis: that DIF signals not only exist but can be effectively detected and exploited to improve model performance. We design two experimental settings: (1) Frozen LLM w/o Align-MLP: We remove the trainable Align-MLP and directly utilize the extracted DIF. This setting investigates whether native LLM activations inherently encode subdomain-aware signals—i.e., whether meaningful associations between user queries and knowledge subdomains can be inferred without explicit alignment. (2) Full DFAMS (w/ Align-MLP): We enable the trainable Align-MLP to utilize DIF to assesse whether modeling the knowledge base on DIF leads to improved alignment and enhanced downstream performance.
@@
We conduct experiments on both the Wiki and Med datasets. Table 4 shows that in setting (1), DFAMS (2000-dim) outperforms both random (2000-dim) and full-layer (3584-dim) baselines, achieving +14.6% accuracy gain on Wiki and +11.78% on Med over the random baseline. These results validate our hypothesis: DIF captures query-subdomain associations and can be directly leveraged, even without further training. Moreover, DFAMS also surpasses the Full baseline, yielding +3.69% and +2.01% accuracy improvements on the Wiki and Med datasets, respectively. In setting (2), with Align-MLP enabled, DFAMS achieves +3.54% and +2.61% higher accuracy than the random and full-layer baselines on Wiki, and +5.34% and +3.45% on Med. These results highlight the benefit of leveraging DIF for knowledge base modeling, resulting in better alignment and improved downstream performance. Additional results and analyses are provided in Appendix H.
@@ 4.4 Sensitivity Analysis (RQ3)
As shown in Figure 3 (right), when top-$T$ is 3, it achieves the best balance (84.2% accuracy, 53.8% recall). Lower $T$ values miss relevant sources, while higher $T$ dilutes document allocation across knowledge bases, reducing key information retrieval.
5 Conclusion
We propose DFAMS, a novel FR framework that explicitly models the DIF within LLMs to enhance query understanding and cross-source knowledge alignment. By leveraging gradient-based neuron attribution and Shapley value estimation, DFAMS identifies latent neural activation paths that reflect user intent and subdomain relevance. The framework also introduces a Multi-Prototype Knowledge Alignment and Routing strategy, which enables fine-grained modeling of individual knowledge bases. Our experiments show that DFAMS consistently outperforms existing FR methods , confirming its effectiveness in resolving semantic ambiguities and improving cross-source routing in complex settings.
Limitations
While DFAMS achieves consistent improvements in federated retrieval, several limitations still remain. First of all, the framework relies on extracting Dynamic Information Flow (DIF) signals from large transformer-based LLMs; while we demonstrate scalability to smaller models, the applicability of DIF-based modeling to fundamentally different architectures has not been thoroughly examined. Second, our evaluation is conducted on curated benchmarks with static knowledge bases. In real-world applications, knowledge sources often evolve continuously; when significant updates occur, the framework may require retraining or integration with continual learning strategies, which has not yet been fully explored. Addressing these aspects could further enhance the robustness and applicability of DFAMS in broader federated retrieval scenarios.
@@ 5 Conclusion
Ethical considerations
To evaluate the efficacy of our work, we conducted experiments using five datasets: Wiki, Med, PEP, MMLU, and MIRAGE. Except for PEP, all datasets are publicly available and used in accordance with their respective licenses and terms of use. PEP was obtained and used with proper authorization. The datasets do not contain personally identifiable information, and no human or animal subjects were directly involved in this research.
Appendix A Notations Table
@@ II Number of isolated knowledge bases (KBs)
$\mathcal{D}=\{\mathcal{K}_{1},\dots,\mathcal{K}_{I}\}$ KBs representing non-parameterized knowledge
$R$ Retrieved subset of KBs for answering query
@@
$y$ Ground-truth answer to query $x$
@@
$g_{\text{align}}$ Projection function mapping $\mathbf{z}$ to aligned query embedding $\mathbf{r}$ in the semantic space
@@ Appendix B Algorithm
16:Select top layers $\mathcal{L}_{\text{top}}$ with highest total Shapley mass
17:For each layer $\ell\in\mathcal{L}_{\text{top}}$, select top neuron groups $\mathcal{G}_{\ell}$ of size $G$
18:Construct $\mathcal{N}=\{\,h_{\ell}^{(g)}\mid\ell\in\mathcal{L}_{\text{top}},\,g\in\mathcal{G}_{\ell}\,\}$
19:return $\mathcal{N}$
B.1 Neuron Probing for DIF Extraction
To identify domain-sensitive neurons within pretrained LLMs, Algorithm 1 introduces how to locate the subset most relevant for DIF. The process initiates by iterating over a probing dataset, capturing the model’s loss and computing per-neuron gradients for each data sample. To measure each neuron’s importance regarding domain sensitivity, Shapley values are approximated for all neurons across the LLM’s layers. This approach quantifies each neuron’s marginal contribution to the model’s ability to capture domain-specific information. The cumulative Shapley scores are averaged over all samples in the probing set, after which the algorithm selects the layers with the highest aggregate Shapley mass. Within these layers, the most influential neuron groups are further identified based on group-wise Shapley values. The final output is a compact set of domain-relevant neurons, $\mathcal{N}$, which serve as the basis for extracting DIF representations in subsequent stages.
@@ B.1 Neuron Probing for DIF Extraction
3:Stage 0: DIF Embedding Extraction
@@
B.2 Two-Stage Contrastive Alignment of DIF
Algorithm 2 delineates a two-stage contrastive training procedure designed to optimize DIF embeddings for robust alignment across disparate knowledge bases (KBs). In the first stage, DIF embeddings are extracted for each training instance by probing the pretrained LLM at the identified neurons, yielding a compact representation that encapsulates domain-specific traits. The initial phase, inter-KB alignment, leverages contrastive learning to encourage DIF embeddings from semantically similar content across different KBs to occupy proximate regions in the embedding space. This is achieved by repeatedly updating the alignment network $g_{align}$ to minimize the contrastive loss $\mathcal{L}_{cl}$, drawing together positive pairs and repelling negatives. The second stage, intra-KB alignment, further refines the DIF space by introducing prototype representations for subdomains or classes within each KB, initialized via unsupervised clustering. Here, a prototype-based contrastive loss $\mathcal{L}_{pcl}$ is incorporated alongside the global loss, jointly optimizing $g_{align}$ to produce discriminative representations that not only bridge domains but respect intra-domain structure. The training concludes with the computation of updated prototypes representative of key semantic clusters, returning both the optimized alignment network and the learned prototypes for later inference.
B.3 Adaptive Prototype-Guided Routing
Algorithm 3 proposes a dynamic routing strategy for inference, leveraging the previously learned prototypes to efficiently allocate retrieval resources based on the semantic relevance of incoming queries. Upon receiving a query, the system first computes its DIF embedding by probing the pretrained LLM at the selected neuron set, followed by transformation via the trained alignment encoder. The query embedding is then compared to all domain prototypes using a similarity metric, yielding a relevance score for each prototype. If the highest similarity score falls below a predefined threshold, the system abstains from retrieval, indicating insufficient semantic alignment. Otherwise, the algorithm selects the top-N most relevant prototypes and aggregates similarity scores by knowledge base, proportionally allocating retrieval slots according to their relative relevance. This prototype-guided mechanism enables adaptive, fine-grained routing decisions that efficiently direct retrieval attention to the most promising knowledge sources, facilitating both high precision and scalable inference in multi-domain settings.
@@ B.3 Adaptive Prototype-Guided Routing
Appendix C Retrieval Pipeline and Indexing Strategies
We adopt FAISS Douze et al. (2024) for dense vector indexing across all retrieval settings. For the Wikipedia (Wiki) knowledge base, we follow the clustering strategy in RopMura Wu et al. (2025), partitioning 1M English passages into 10 semantically coherent knowledge bases. Passages are first embedded using Qwen-embedding-v2 Bai et al. (2023) for clustering, and subsequently indexed with all-MiniLM-L6-v2 Wang et al. (2020). For the Medical (Med) knowledge base Zhao et al. (2025), we replicate the knowledge base construction from RAGRoute Guerraoui et al. (2025), creating four distinct sources: PubMed, StatPearls, medical textbooks, and medical Wikipedia, each indexed using all-ßMiniLM-L6-v2. For the Private Enterprise Policy (PEP) knowledge base, which contains internal Chinese-language company documents spanning four sub-knowledge bases, we utilize the GTE-base-zh encoder Li et al. (2023) fr indexing. For each query, DFAMS retrieves the top-10 documents from the dynamically selected source(s).
Appendix D Dataset Construction and Sampling Strategies
Wiki Dataset Construction.
Med Dataset Construction.
Following a similar processing pipeline as Wiki, we construct the training and test sets for the medical domain. The training set includes only “no retrieval” or “single-knowledge-base single-segment” queries to reduce annotation costs. The test set additionally incorporates “cross-knowledge-base multi-segment” and “same-knowledge-base multi-segment” queries to evaluate the system’s generalization ability in realistic multi-knowledge-base collaboration scenarios. The test set contains 2975samples, with 438 requiring cross-knowledge-base multi-document collaboration and 823 requiring same-knowledge-base multi-document collaboration, to assess robustness in real multi-knowledge-base environments.
PEP Dataset Construction.
OOD Dataset Construction.
Following the evaluation paradigm of RAGRoute Guerraoui et al. (2025), we construct lightweight OOD test sets by extracting sub-questions from MMLU Hendrycks et al. (2020) and MIRAGE Xiong et al. (2024) that are most relevant to the topics covered by the existing four knowledge bases. For MMLU, we retain 1,222 questions that are potentially related to the WIKI knowledge base; for MIRAGE, we filter 1,546 open-domain QA samples with the highest entity co-occurrence with the four knowledge bases. Both subsets lack corresponding golden-label documents and are used solely to evaluate the model’s robustness and knowledge generalization under domain shift and non-retrieval conditions.
Appendix E Baseline Implementation Details
No-RAG.
Merged-RAG.
Prompt.
CoT Prompt.
A knowledge bases selection where a 70B teacher model is used to perform corpus routing, but with chain-of-thought (CoT) prompting Wei et al. (2022). Compared to Prompt, this variant enhances reasoning by explicitly incorporating intermediate steps during knowledge bases classification and downstream answer generation.
SFT.
A supervised fine-tuning baseline trained on the $D_{train}$ dataset using a cross-entropy loss. We fine-tuned two base models, Qwen2.5-7B and LLaMA3.1-8B, to predict the correct knowledge base for each query. The training specifically targets improving knowledge base selection accuracy.
RopMura.
A recent joint retrieval and routing method Wu et al. (2025). As our focus is on knowledge bases selection, we isolate and evaluate only the knowledge bases selection module. Multi-turn dialog components are disabled for fair comparison.
RAGRoute.
For each knowledge bases, we train an MLP-based router whose architecture and training settings exactly match those of our Align-MLP, ensuring a fair comparison Guerraoui et al. (2025). All methods share the same encoder and retriever: we use all-MiniLM-L6-v2 for Wiki and Med, and gte-base-zh for PEP.
Appendix F Metric Definitions and Evaluation Configuration
We evaluate DFAMS using three complementary metrics:
Cls Acc.
Recall.
QA.
This metric evaluates the final response quality. For multiple-choice questions, we extract the predicted option(s) (e.g., A/B/C/D) from the model output and compare them against the ground-truth answer. For open-ended queries, responses are scored by an LLM-based judge Zheng et al. (2023) that assesses factual correctness and fluency. The score ranges from 0 to 10.
Appendix G Implementation Details
To train the Multi-Prototype Knowledge Alignment module of DFAMS, we adopt a two-stage process. In the first stage, each private knowledge base independently extracts its own DIF representations from its local training data $\mathcal{D}_{\text{train}}$, where the activations of the selected neuron groups across designated layers are concatenated and pooled across tokens (e.g., using AverageToken) to form the final DIF representations. These DIFs capture the semantic characteristics of the local knowledge bases without exposing the underlying textual content, and are securely stored and transferred to the aligner for training. In the second stage, the DIF representations from all participating knowledge bases are aggregated and fed into the Align-MLP for joint optimization. The aligner is trained with an inter-KB supervised contrastive loss and an intra-KB multi-prototype contrastive loss (PCL), with prototypes initialized via KMeans clustering. This two-stage design effectively reduces redundant computations, improves overall training efficiency, and satisfies the data-locality and privacy-preservation requirements inherent to federated retrieval scenarios.
We train the Align-MLP aligner using the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a cosine decay schedule. Unless otherwise specified, the batch size is set to 64, the temperature parameter to 0.07, and the number of contrastive learning (CL) epochs to four, during which only the CL objective is optimized. All models are trained for six epochs in total, with the final one to two epochs jointly optimizing both the CL and PCL. The DIF representations are extracted from the mlp.up_\_proj components of the 26th and 27th transformer layers, where attribution-based analysis identifies neuron groups that are most relevant to the retrieval objective. Specifically, neurons are grouped in sets of 20, and the top 50 groups (corresponding to 1,000 neurons per layer) with the highest attribution scores are selected for DIF construction. The Align-MLP shares the same architecture as the probing network used in Probing-RAG, consisting of three fully connected layers with intermediate SiLU activations, layer normalization, and dropout regularization. The hidden dimension of the MLP is set to 512, and its input and output dimensions match those of the DIF representations to ensure alignment consistency.
We train the Align-MLP aligner using the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a cosine decay schedule. Unless otherwise specified, the batch size is set to 64, the temperature parameter to 0.07, and the number of contrastive learning (CL) epochs to four, during which only the CL objective is optimized. All models are trained for six epochs in total, with the final one to two epochs jointly optimizing both the CL and PCL objectives. The DIF representations are extracted from the mlp.up_\_proj components of the 26th and 27th transformer layers, where attribution-based analysis identifies neuron groups most relevant to the retrieval objective. Specifically, neurons are grouped in sets of 20, and the top 50 groups (corresponding to 1,000 neurons per layer) with the highest attribution scores are selected for DIF construction. The Align-MLP follows the same architecture as the probing network used in Probing-RAG, consisting of three fully connected layers with intermediate SiLU activations, layer normalization, and dropout regularization. The hidden dimension of the MLP is set to 512, and both its input and output dimensions are aligned with the 2,000-dimensional DIF feature space. Across all experiments, the loss weighting parameter is fixed to $\alpha=0.95$, meaning that the PCL term contributes 0.05 of the total loss. Other hyperparameter configurations are summarized in Table 7.
Hyperparameter Wiki Med PEP MMLU MIRAGE
Number of KBs 10 4 4 10 4
Learning Rate 2e-4 1e-4 1e-4 2e-4 1e-4
Epochs 13 20 6 13 20
CL Epochs 12 22 4 12 22
Batch Size 64 64 64 64 64
Threshold 0.85 0.95 0.80 0.85 0.95
Table 7: Hyperparameter summary across different datasets and experimental scenarios.
Appendix H Additional Results on Dynamic Information Flow
To verify our hypothesis regarding the presence of DIF, we analyzed the Shapley value heatmaps across four model sizes, as shown in Figure 4. The results reveal a consistent trend: the Shapley values increase in the early-to-middle layers (e.g., around layers 1–6 in Qwen2.5-7B), decrease in the intermediate layers, and rise again in the deeper layers (e.g., layers 26–27). This pattern suggests that shallow layers primarily capture intent-related signals, while deeper layers activate and integrate domain-specific knowledge. This observation supports our decision to extract DIF representations from the latter layers.
Figure 4: Heatmaps of Aggregated Shapley Values Across Layers and Neuron Groups for Qwen2.5-7B, Qwen2.5-3B, Qwen2.5-5B, and LLaMA3.1-8B Models
[…]
26 7.53 0.671 0.3765 72.70
27 -10.19 1.030 1.1895 72.51
4 3.13 0.662 0.1357 56.61
3 3.89 0.615 0.1339 46.17
15 4.36 0.635 0.0737 27.83
Table 8: Layer-wise attribution results on the PEP dataset.
[…]
Appendix I Effect of Backbone Model Size.
[…]
Qwen2.5-0.5B 81.23 50.75
Qwen2.5-3B 83.91 51.39
Qwen2.5-7B 85.03 53.83
Appendix J Effect of Learning Rate
[…]
Cls Acc ($\uparrow$) 81.51 83.56 85.03 83.22 54.25
Table 10: Cls Acc of different learning rates using DFAMS.
Appendix K Computational Resources and Software Environment
Experiments were conducted on a server equipped with dual Intel Xeon E5-2680 v4 CPUs (56 cores, 112 threads), 8 NVIDIA RTX 3090 GPUs (24GB each), and 377 GB of main memory, running Ubuntu 18.04.6 LTS. Python 3.11.10 was used with PyTorch 2.4.0, and packages were managed via Conda 23.5.2. Model training took approximately 1 minute per epoch on average, depending on the experimental setting. Data preprocessing, including extraction of around 1,000 DIF samples ( 4 MB), required about 10 minutes to obtain the DIF representations. During inference, generating QA accuracy results for 1,000 samples required approximately 5 hours on the same hardware configuration. All models and software packages, including HuggingFace Transformers 4.44.0, SpaCy 3.7.0, NLTK 3.8.1, and related dependencies, were used with default or explicitly stated parameter settings.
Appendix L The Use of Large Language Models
In this work, Large Language Models (LLMs) were used for language polishing and coding assistance. Specifically, LLMs supported refining the clarity and grammar of the manuscript, improving stylistic quality, and suggesting code snippets or troubleshooting strategies. All content generated by LLMs was carefully reviewed and verified by the authors before inclusion. The research design, critical analyses, and all final decisions were independently conducted by the authors. LLMs were not involved in generating new research ideas or conclusions.
Appendix M Prompt